Calculate the isoelectric point of glutamic acid. How does the isoelectric point of glutamic acid differ from glycine.
1
Expert's answer
2020-12-14T05:35:01-0500

pI=(pKa1+pKa3)÷2

pKa1=2.19

pKa2=9.67

pKa3=4.25


pI=(2.19+4.25)÷2

=6.44÷2

=3.22

The pI of glycine is 5.97. It is different from glutamic acid because glycine contains a hydrogen as its side chain instead of a carbon as is the case with glutamic acid.